# Concurrent Composition for Differentially Private Continual Mechanisms

### v2

@@ 7. Application (Continual Monotone Histogram Query Problem)
We next show an application of our technique to the continual monotone histogram query problem. The continual monotone histogram query problem is defined as follows.
@@
Let $d\in\mathbb{N}$ and let $q:\mathbb{N}^{d}\rightarrow\mathbb{N}$ be a function that is monotone in its input. The input is a finite stream of row vectors $x^{1},\dots,x^{T}$ with $x^{t}\in\{0,1\}^{d}$ for all $t\in[T]$. Let $\sim$ be a neighbor relation on ($\left(\{0,1\}^{d}\right)^{*}$ where two input streams are neighboring if they are identical except for a single step. The output of the (differentially private) continual monotone histogram query problem at each time step $t$ is (an additive approximation to) $q(t)=q((\sum_{\ell=1}^{t}x_{i}^{\ell})_{i\in[d]})$.
Next we define query sensitivity.
@@
The function $q:\{0,1\}^{d}\rightarrow\mathbb{R}$ is $1$-sensitive if $\max_{x\sim y}|q(x)-q(y)|$ is at most $1$.
Henzinger, Sricharan, and Steiner [12] presented the best known differentially private mechanism for the monotone histogram query problem, where where the query is $1$-sensitive. For $\epsilon$-differential privacy it has, with probability at least $1-\beta$, an additive error of $O((d\log^{2}(dq^{*}/\beta)+\log T)\epsilon^{-1})$, and for $(\epsilon,\delta)$-differential privacy, it has an additive error of $O((d\log^{1.5}(dq^{*}/\beta)+\log T)\epsilon^{-1}\log\frac{1}{\delta})$, where $q^{*}=\max_{t\in[T]}q(t)$. The privacy of the mechanism is proven “from scratch” using the concept of a privacy game [13] and the latter bound only holds against an oblivious adversary.
@@
[…]
@@
procedure SVT-call($x,\texttt{Thresh}$):
@@
Let $f_{\sim}$ be a verification function that given a message sequence $m_{1},\dots,m_{t}$, returns $\top$ if and only if for each $j\in[t]$, the message $m_{j}$ is of the form $(m_{j}^{0},m_{j}^{1})$ and the message sequences $\sigma_{0}=(m_{1}^{0},\dots,m_{t}^{0})$ and $\sigma_{1}=(m_{1}^{1},\dots,m_{t}^{1})$ satisfy $\sigma_{0}\sim\sigma_{1}$. As part of their privacy proof, the following is shown in [12], which we restate in our own language.
@@
If $q$ is a $1$-sensitive function, then for every $\epsilon$ and $h$, the mechanism $\textrm{SVT}_{\epsilon/6,q,h}$ is $\epsilon/3$-DP w.r.t. $f_{\sim}$.
For positive integers $t$ and $j$, $\beta\in(0,1]$, and $\epsilon>0$, let $\gamma(t,j,\beta,\epsilon)$ and $\xi(t,j,\beta,\epsilon)$ are some deterministic functions. Given an input stream $x_{1},x_{2},\dots$ with $x_{i}\in\{0,1\}^{d}$, the algorithm partitions the input stream into subsequences, called intervals. At the end of each interval, it gives the $d$-dimensional sum of all the inputs of the interval to a differentially-private histogram mechanism H. Note that H can be used directly to solve the problem; however, the new mechanism HSS has better accuracy bounds in many cases.
[…]
@@
[…]
@@
for $t\in\mathbb{N}$ do
     $c_{i}\leftarrow c_{i}+x^{(t)}_{i}$ for all $i\in[d]$,
[…]
         $h=\texttt{H-add}(c)$
@@
Algorithm 4 describes a CM HSS with query set $\{0,1\}^{d}$ and answer set $\mathbb{N}^{d}$. The state of HSS consists of the variables $(h,c,j,out,\texttt{Thresh})$, initialized with $(0^{d},0^{d},1,q(0^{d}),\gamma(1,1,\beta,\epsilon))$. The summation $q(c+h)+\mathrm{Lap}(3/\epsilon)$ in this algorithm can be rewritten as an in instance of Laplace mechanism with privacy parameter $\epsilon/3$ and input $q(c+h)$. Consequently, HSS executes one instance of a histogram mechanism, H, an unbounded number of SVT instances, and an unbounded number of Laplace mechanisms. The Laplace mechanism with privacy parameter $\epsilon/3$ is a NIM that is known to be $\epsilon/3$-DP when receiving real values differing by at most $1$. In other words, the Laplace mechanism is a CM that is $\epsilon/3$-DP w.r.t. a verification function $f_{\mathrm{Lap}}$ that returns $\top$ if and only if it receives a single input message $m=(m_{0},m_{1})\in\mathbb{R}^{2}$ satisfying $|m_{0}-m_{1}|\leq 1$.
@@
The output of H is a vector in $\mathbb{R}^{d}$, while the output of the Laplace mechanism is a scalar in $\mathbb{R}$. Rounding up each real-valued output to the nearest integer does not affect privacy guarantees (by the post-processing lemma) but increases the additive error of each mechanism by at most $1$. This extra error is negligible compared to their original error. Therefore, we assume that both H and the Laplace mechanism have discrete answer distributions. Since the answer set of SVT is finite, it also has discrete answer distributions.
By design, HSS uses the inputs $x^{(t)}$ only to determine the inputs of its sub-mechanisms, and every other action is taken solely based on the privatized outputs of the sub-mechanisms. This allows us to apply Theorem 6.1 to analyze the privacy of HSS based on the privacy parameters of its sub-mechanisms.
By Lemma 7.3, each instance of SVT created by HSS is $\epsilon/3$-DP. Furthermore, for any two neighboring input streams, fixing the history of privatized outputs of the sub-mechanisms, HSS provides identical inputs to all but one SVT instance, which receives neighboring inputs. By Theorem 5.2, the parallel concurrent composition of all SVT instances is $\epsilon/3$-DP.
Similarly, HSS gives identical real-valued inputs to all but one instance of the Laplace mechanism, where the differing inputs vary by at most $1$. This implies that the concurrent composition of all Laplace mechanisms is also $\epsilon/3$-DP. Each concurrent composition is itself a continual mechanism. Additionally, it is given that H is $\epsilon/3$-DP. By basic composition, the composition of three instances of $\mathrm{RR}_{\epsilon/3,0}$ is $\epsilon$-DP. Therefore, by Theorem 4.10, the concurrent composition of all mechanisms—H, SVT instances, and Laplace mechanisms—is $\epsilon$-DP. Applying Theorem 6.1, the mechanism HSS is $\epsilon$-DP.
In the case where H is $(\epsilon/3,\delta)$-DP, since the composition of $\mathrm{RR}_{\epsilon/3,0}$, $\mathrm{RR}_{\epsilon/3,0}$, and $\mathrm{RR}_{\epsilon/3,\delta}$ is $(\epsilon,\delta)$-DP, the mechanism HSS is also $(\epsilon,\delta)$-DP. Both privacy guarantees provided in this section hold against adaptive adversaries.
Acknowledgments
Appendix A Overview and Modification of [18]
In this section, we overview and modify the results of [18]. The original proof relies on the existence of a constant $T$ that upper bounds the number of queries adversaries ask; however, our modified construction does not. Moreover, an algorithmic construction in [18]’s proof needs the answer set to be finite. Using mathematical tools, we formalize their construction and weaken this assumption to having discrete answer distributions. We show:
Lemma A.1.

### v3

@@ 7. Application (Continual Monotone Histogram Query Problem)
We next show an application of our technique to the continual monotone histogram query problem. So far, we defined differential privacy (DP) with respect to (w.r.t.) verification functions. To be consistent with the literature, in this section we use the notion of DP w.r.t. “neighboring input sequences” instead:
Remark 7.1.
Let $\sim$ be a binary relation on a family of message sequences. Throughout this section, we say that a mechanism is $(\epsilon,\delta)$-DP w.r.t. the neighbor relation $\sim$ (or simply it is $(\epsilon,\delta)$-DP) if the mechanism is $(\epsilon,\delta)$-DP w.r.t. to the verification function $f_{\sim}$ against adaptive adversaries, where $f_{\sim}$ is defined as follows: For every message sequence $m_{1},\dots,m_{t}$, $f_{\sim}(m_{1}\dots,m_{t})=\top$ if and only if the message sequence is suitable (implying $m_{j}=(m_{j}^{0},m_{j}^{1})$ for every $j\in\{1,\dots,t\}$) and the sequences $\sigma_{0}=(m_{1}^{0},\dots,m_{t}^{0})$ and $\sigma_{1}=(m_{1}^{1},\dots,m_{t}^{1})$ satisfy $\sigma_{0}\sim\sigma_{1}$.
@@
Let $d,T\in\mathbb{N}$ and let $q:\mathbb{N}^{d}\to\mathbb{N}$ be a function that is monotone in its input. In the (differentially private) continual monotone histogram query problem, a continual mechanism $\mathcal{M}$ receives $T$ inputs of row vectors $x^{1},\dots,x^{T}$, each in $\{0,1\}^{d}$. Define $\sim$ to be a neighbor relation on $\left(\{0,1\}^{d}\right)^{*}$ where two input streams are neighboring if they are identical except for a single step. Differential privacy for $\mathcal{M}$ is defined with respect to this neighbor relation (see Remark 7.1). The output of $\mathcal{M}$ at each time step $t$ is (an additive approximation to) $q(t)=q((\sum_{\ell=1}^{t}x_{i}^{\ell})_{i\in[d]})$.
@@
Recall function $q:\{0,1\}^{d}\to\mathbb{R}$ and neighbor relation $\sim$ from Definition 7.2. $q$ is said to be $1$-sensitive if $\max_{x^{0}\sim x^{1}}|q(x^{0})-q(x^{1})|$ is at most $1$.
Henzinger, Sricharan, and Steiner [13] presented the best known differentially private mechanism for the monotone histogram query problem, where the query is $1$-sensitive. For $\epsilon$-differential privacy it has, with probability at least $1-\beta$, an additive error of $O((d\log^{2}(dq^{*}/\beta)+\log T)\epsilon^{-1})$, and for $(\epsilon,\delta)$-differential privacy, it has an additive error of $O((d\log^{1.5}(dq^{*}/\beta)+\log T)\epsilon^{-1}\log\frac{1}{\delta})$, where $q^{*}=\max_{t\in[T]}q(t)$. The privacy of the mechanism is proven “from scratch” using the concept of a privacy game [14] and the latter bound only holds against an oblivious adversary.
@@
The mechanism HSS, described in Algorithm 4, employs a variant of the Sparse Vector Technique mechanism, SVT, and a $d$-dimensional counter mechanism, d-counter, as submechanisms. We start by defining these mechanisms and then explain HSS.
Sparse vector technique mechanism (SVT). The mechanism SVT is described in Algorithm 3 using two operations. The first, $\texttt{SVT-init}(\epsilon,q,h)$, initializes the mechanism with a privacy parameter $\epsilon$, a query $q$, and an initial $d$-dimensional vector $h$. This procedure indeed describes the creation of a continual mechanism $\textrm{SVT}_{h,\epsilon,q}$. The second operation, $\texttt{SVT-update}(x,\texttt{Thresh})$, maps inputs $(x,\texttt{Thresh})\in\{0,1\}^{d}\times\mathbb{R}$ to outputs in $\{\top,\bot\}$. The mechanism maintains a running sum of all inputs $x$ and, in a differentially private manner, checks whether applying $q$ to the running sum yields a value at least Thresh. If so, it outputs $\top$ and halts. Otherwise, it outputs $\bot$ and continues. These two procedures are given in detail in Algorithm 3. The command $\mathrm{Lap}(b)$ denotes a random sample from the Laplace distribution with scale parameter $b$.
To clarify the connection between these pseudocodes and our definition of continual mechanisms (Definition 3.1), we note that the state of $\textrm{SVT}_{h,\epsilon,q}$ is the tuple of its variables. When $\texttt{SVT-init}(\epsilon,q,h)$ is invoked, the continual mechanism $\textrm{SVT}_{h,\epsilon,q}$ with the single initial state $(\epsilon,q,h,\tau=\text{null})$ is created. In the pseudocode of SVT-init, the variable $\tau$ takes value. In our setting, such initialization is considered to happen when $\textrm{SVT}_{h,\epsilon,q}$ receives its first input. The SVT-update operation then describes how $\textrm{SVT}_{h,\epsilon,q}$ maps its current state (i.e., set of variables) and an input message $m=(x,\texttt{Thresh})\in\{0,1\}^{d}\times\mathbb{R}$ to a new states (with updated variables) and a response $m^{\prime}\in\{\top,\bot\}$.
@@
procedure SVT-update($x,\texttt{Thresh}$):
@@
As part of their privacy proof, the following is shown in [13]:
@@
If $q$ is a $1$-sensitive function, then for every $\epsilon$ and $h$, the mechanism $\textrm{SVT}_{\epsilon/6,q,h}$ is $\epsilon/3$-DP w.r.t. the neighbor relation $\sim$ in Definition 7.2.
$d$-dimensional continual counter. The mechanism d-counter is described by two operations. The first, $\texttt{d-counter-init}(\epsilon)$, creates $d$ differentially private continual binary counters with privacy parameter $\epsilon/d$. The second, $\texttt{d-counter-update}(x)$, gives the elements of its $d$-dimensional input $x$ to their corresponding binary counters and returns the aggregated outputs as a $d$-dimensional vector. We denote this continual mechanism by $\texttt{d-counter}_{\epsilon}$. In another version, $\texttt{d-counter-init}(\epsilon,\delta)$ creates $d$ differentially private continual binary counters with privacy parameters $\epsilon/d$ and $\delta/d$. We represent this continual mechanism by $\texttt{d-counter}_{\epsilon,\delta}$.
By definition, mechanisms $\texttt{d-counter}_{\epsilon}$ and $\texttt{d-counter}_{\epsilon,\delta}$ are the concurrent composition of $d$ many $\epsilon/d$- and $(\epsilon/d,\delta/d)$-DP fixed private continual counters. We can assume that these continual counters return an integer value at each point of time, and thus they have discrete answer distributions. If the output space is real, we use a counter with integer outputs that post-processes the original counter by rounding its output up to the nearest integer. By the post-processing lemma, the privacy guarantees still hold and the additive error is only negligibly affected by at most $1$. Therefore, Theorem 4.9 is applicable and by the basic composition, we have:
Corollary 7.5.
Mechanisms $\texttt{d-counter}_{\epsilon}$ and $\texttt{d-counter}_{\epsilon,\delta}$ are $\epsilon$-DP and $(\epsilon,\delta)$-DP, respectively.
Continual histogram query mechanism. The mechanism HSS, described in Algorithm 4, is initialized with a privacy parameter $\epsilon>0$ and an accuracy parameter $0<\beta\leq 1$ and receives a vector $x\in\{0,1\}^{d}$ as input at each time step. This mechanism partitions the input stream into subsequences, called intervals. At the end of each interval, it forwards the $d$-dimensional sum of all the inputs of the interval to the differentially private $d$-dimensional continual counter d-counter. (Note that d-counter can be used directly to solve the problem; however, HSS achieves better accuracy bounds in many cases.) The mechanism HSS maintains two $d$-dimensional vectors: (1) variable $c$, where $c_{i}$ stores the exact sum of coordinate $i$ within the current interval, and (2) variable $h$, where $h_{i}$ stores the approximate sum of coordinate $i$ over all previous intervals computed by d-counter. Both vectors are initialized to $0^{d}$.
At the beginning of each interval (including the initial time step), HSS creates a fresh instance of the Sparse Vector Technique mechanism SVT with privacy parameter $\epsilon/6$, query $q$, and initial vector $h=0^{d}$ (denoted by $\textrm{SVT}_{\epsilon/6,q,h}$). The mechanism discards the previous instance. HSS keeps track of the current time step and the current interval using variables $t$ and $j$, respectively. In Algorithm 4, $\gamma$ and $\xi$ are some deterministic functions, mapping $t,j,\beta,\epsilon$ to a real value.
The mechanism HSS also stores a threshold variable Thresh, initially set to $\gamma(1,1,\beta,\epsilon)$. Upon receiving an input $x$ at step $t$, the mechanism updates $c$ and invokes $\texttt{SVT-update}(x,\texttt{Thresh})$. If the response is $\top$, the current interval ends: $c$ is passed to d-counter, $h$ is updated with the output of d-counter, and $c$ is reset to $0^{d}$. Since SVT halts after returning $\top$, a new instance $\textrm{SVT}_{\epsilon/6,q,h}$ is created. In this case, HSS privately evaluates $q(c+h)$ via the Laplace noise and compares the noisy result to $\texttt{Thresh}-\xi(t,j,\beta,\epsilon)$. If the inequality holds, Thresh is suitably increased. Moreover, independent of the outcome of this comparison, Thresh is updated again to reflect the increase in the number of intervals. Then, regardless of the output of $\texttt{SVT-update}(x,\texttt{Thresh})$, HSS updates Thresh once more to account for the increased number of time steps. Finally, HSS outputs $h$.
@@
procedure Hist-init($\epsilon,q,\beta$):
  $\texttt{d-counter-init}(\epsilon/3)$ (or $\texttt{d-counter-init}(\epsilon/3,\delta)$)
@@
  $t\leftarrow 0$
end procedure
procedure Hist-update($x$):
  $t\leftarrow t+1$
  $c_{i}\leftarrow c_{i}+x_{i}$ for all $i\in[d]$,
  if $\texttt{SVT-update}(x,\texttt{Thresh})==\top$ then
   $h=\texttt{d-counter-update}(c)$
@@
In Algorithm 4, the summation $q(c+h)+\mathrm{Lap}(3/\epsilon)$ can be rewritten as an instance of Laplace mechanism with privacy parameter $\epsilon/3$ and input $q(c+h)$. Consequently, the mechanism HSS executes:
  1. (1)
    one instance of the $d$-dimensional counter d-counter, which consists of $d$ instances of a scalar counter,
  2. (2)
    an arbitrary number of SVT instances, and
  3. (3)
    an arbitrary number of Laplace mechanisms.
By Corollary 7.5, d-counter is $\epsilon/3$-DP (or $(\epsilon/3,\delta)$-DP when approximate counters are used). By Lemma 7.4, each SVT instance is $\epsilon/3$-DP. Finally, the Laplace mechanism with parameter $\epsilon/3$ is a non-interactive mechanism (NIM) that is well known to satisfy $\epsilon/3$-DP on neighboring real inputs differing by at most $1$. The idea is to show that the concurrent composition of multiple instances of these mechanisms created by HSS is $\epsilon$-DP (or $(\epsilon,\delta)$-DP when d-counter is approx DP). Then, we will then show that HSS satisfies the required properties in Corollary 6.5, and thus it has the same privacy guarantees as the concurrent composition. As concurrent compositions theorems in this paper require the composed mechanisms to have discrete answer distributions, we first show that we can assume the output space of d-counter and Laplace mechanism is discrete. Since SVT has a finite output space, its answer distributions are also discrete.
@@
Discretization of outputs. As argued earlier, the counters executed by d-counter outputs integer values, and thus d-counter has discrete answer distributions. Likewise, we can assume that the Laplace mechanism in HSS rounds its noisy real-valued output to the nearest integer. By the post-processing lemma, this rounding does not degrade privacy, and by definition, it increases the additive error by at most $1$, which is unimportant for the accuracy guarantees of [13]. Hence, both d-counter and the Laplace mechanism can be treated as having discrete output distributions.
Pseudocode privacy analysis. Recall from Section 6.3 that decision variables of a mechanism are either (i) outputs of submechanisms, or (ii) variables computed solely as functions of other decision variables. In Algorithm 4, the variables $h$, outout, $j$, $t$, and Thresh are decision variables, while $c$ is not: Variable $h$ denotes the privatized outcomes of d-counter. Variable outout equals $q(h)$, where $h$ is a decision variable. Index $j$ is initially $1$ and is increased by $1$ when the privatized output of SVT is $\top$. Time step $t$ is initially 0 and is increase by $1$ at each step, independent of the input and other variables. Finally, variable Thresh is initialized and updated to values determined by $\epsilon,\beta,j,t$, where $\epsilon,\beta$ are fixed and $j,t$ are decision variables. It is important to note that the decision of updating variables $h$ (and consequently outout) and Thresh depends only on the privatized output of SVT.
We first show that the destination property holds. For that we have to show that the decision of whether to produce an output, create a new submechanism (and if so, which one) or interact with an existing one (and if so, which one) are solely based on decision variables. All submechanisms of HSS are instantiated either in the initial step (and, thus, whether it receives input is independent of non-decision variables and the inputs $x$) or based on outputs of the private submechanism SVT. The active SVT instance receives input at every step, and thus whether it receives input is independent of non-decision variables and the inputs $x$. Also, whether the mechanism d-counter gets updated depends only on the privatized output of SVT. Finally, HSS returns an output after the final update of Thresh, without any condition on the inputs $x$ or the non-decision variables. Thus, the destination property in Corollary 6.5 holds. Moreover, the final output of HSS is outout, which is a decision variable, so the response property in that corollary also holds. It remains to show HSS satisfies the $f\to f^{\prime}$ property, where $f$ is the verification function corresponding to the neighbor relation $\sim$ in Definition 7.2 and $f^{\prime}$ is a verification satisfying that the $f^{\prime}$-concurrent composition of the continual mechanisms is $(\epsilon,\delta)$-DP.
Fix some values for all decision variables at every time step. Conditioned on these values, when HSS processes two neighboring input sequences, the resulting inputs to its submechanisms behave as follows:
  • d-counter receives neighboring sequences,
  • all but one instance of SVT receive identical inputs, with at most one receiving neighboring inputs, and
  • all but one Laplace mechanism receive identical inputs, with at most one receiving neighboring inputs.
Hence, the concurrent composition of all submechanisms of HSS can be decomposed into three concurrent compositions: (1) all Laplace mechanisms, (2) all SVT instances, and (3) the $d$ continual counters forming d-counter. The first two are the parallel compositions of purely differentially private mechanisms. By Corollary 5.5, both are $\epsilon/3$-DP. The third one is, by construction, the concurrent composition of $d$ fixed continual counters. In Corollary 7.5, we showed that this composition, i.e., d-counter, is $\epsilon/3$-DP (or $(\epsilon/3,\delta)$-DP when approx DP counters are used).
By basic composition and Theorem 4.9, the overall concurrent composition of all three groups of mechanisms is $\epsilon$-DP when d-counter uses pure counters, and $(\epsilon,\delta)$-DP when approximate counters are used. (More precisely, it is DP w.r.t. a verification function $f^{\prime}$ that ensures all submechanisms but one SVT instance, one Laplace mechanism, and the $d$ counters of d-counter are assigned pairs of identical inputs, while the first and second inputs for each of these mechanisms form neighboring sequences.) We note that the concurrent compositions are themselves continual mechanisms and can be composed; an explicit example of composing two compositions in the message-passing model is given in the proof of Lemma 5.11.
By Corollary 6.5, since HSS satisfies the destination, response, and neighboring-mapping properties, and since the concurrent composition of its submechanisms is $\epsilon$-DP (or $(\epsilon,\delta)$-DP), the mechanism HSS satisfies the same privacy guarantee. We recall that every privacy guarantees in this section holds against adaptive adversaries.
8. Reduction to Interactive Randomized Response
Lyu [19] shows that for every $(\epsilon,\delta)$-DP IM $\mathcal{M}$ and every pair of neighboring initial states $s_{0},s_{1}$ for $\mathcal{M}$, there exists an IPM $\mathcal{P}$ such that for each $b\in\{0,1\}$, the mechanisms $\mathcal{M}(s_{b})$ and $\mathcal{P}\!\circ^{*}\!\mathrm{RR}_{\epsilon,\delta}(b)$ are equivalent. Meaning that, $\mathcal{P}$ can simulate $\mathcal{M}(s_{b})$ given access only to the output of a randomized response mechanism $\mathrm{RR}_{\epsilon,\delta}(b)$. In this section, inspired by this work, we construct a new IPM $\mathcal{P}$ that, rather than simulating $\mathcal{M}$ based on the outcome of a single instance of $\mathrm{RR}_{\epsilon,\delta}$, operates on the output of $\mathrm{RR}_{0,\delta}$ and only when necessary uses the output of $\mathrm{RR}_{\epsilon,0}$. In other words, instead of interacting with the NIM $\mathrm{RR}_{\epsilon,\delta}$, $\mathcal{P}$ interacts with the IM $\mathrm{IRR}_{\epsilon,\delta}$, and under certain conditions, it queries $\mathrm{IRR}_{\epsilon,\delta}$ at most once.
In addition, our construction removes or relaxes the following implicit assumptions in Lyu’s proof:
  • Lyu’s result relies on the existence of a constant $T$ that upper bounds the number of queries adversaries ask. This restriction comes from a backward definition of a so-called control function. We define a function which is the limit over this control function and, using monotone convergence, show that this “limit function” satisfies all the necessary conditions to be used in the rest of Lyu’s proof. This extends the result to an unbounded number of rounds of interaction.
  • Lyu’s proof uses an algorithmic construction iterating over the answer set, implying that the answer set must be finite. Instead with a mathematical approach, we extend the proof to work with a countably infinite answer set.
Our result is formalized in Lemma 5.6, which is restated below:
Lemma (Restatement of Lemma 5.6).